# Baker–Akhiezer specialisation of joint eigenfunctions for hyperbolic relativistic Calogero–Moser Hamiltonians

### v1

@@ 1 Introduction
where $(T_{q,x_{l}}f)(x_{1},\ldots,x_{k},\ldots,x_{N})=f(x_{1},\ldots,qx_{l},\ldots,x_{N})$. (The precise relationship can be gleaned from (2.6) and (3.4).) They act on the space of symmetric polynomials in $N$ variables, on which they are simultaneously diagonalised by the symmetric $GL_{N}$ type Macdonald polynomials; see e.g. [Mac95, Sto21]. For parameter values of the form $t=q^{m}$, $m\in{\mathbb{Z}}$, Etingof and Styrkas [ES98] and Chalykh [Cha02] constructed and studied non-symmetric joint eigenfunctions of Baker–Akhiezer (BA) type and obtained, in particular, a generalized Weyl character formula for the Macdonald polynomials, first conjectured by Felder and Varchenko [FV97], where the BA-function replaces the exponential function.
@@
For the coupling values $g=m\hbar$, $m\in\mathbb{Z}_{+}$, under consideration, our main result provides an expansion of symmetric joint eigenfunctions in terms of eigenfunctions that are ‘asymptotically free’ deep inside Weyl chambers $x_{\sigma(1)}>x_{\sigma(2)}>\cdots>x_{\sigma(N)}$, $\sigma\in S_{N}$. In the case of the Macdonald operators and generic parameter values, such asymptotically free solutions, often referred to as (q-)Harisch–Chandra series, have been studied in detail by Letzter and Stokman [LS08], van Meer and Stokman [vMS10] as well as Noumi and Shiraishi [NS12]; and Stokman [Sto14] derived a corresponding ($c$-function) expansion of Cherednik’s basic hypergeometric function associated to root systems. To obtain an analogous expansion for the joint $S_{r}$-eigenfunction from [HR14] in the case of generic $g$-values is, I believe, an interesting open problem to which I hope to return elsewhere. In the $N=2$ case, the first steps in this direction can be found in a recent paper by di Francesco et al. [FKKSS24], in which they obtained a number of interesting results on these eigenfunction, including how Harish–Chandra series are recovered by suitably closing contours in one of their integral representations and computing residues.
The remainder of the paper is structured as follows: In Section 2, we briefly recall definitions and properties pertaining to joint eigenfunctions of the Ruijsenaars operators (1.1) and BA-(eigen)functions for the Macdonald operators (1.3); and, in Section 3, we give the precise formulations and proofs of our results.
@@ 2.1 Joint eigenfunctions
see loc. cit. Prop. III.1 and Prop. III.2. Indeed, using these difference equations, it is readily verified that
$\begin{split}A_{r,\delta}(x)&\equiv W_{N}(x)^{-1/2}H_{r,\delta}(b;x)W_{N}(x)^{1/2}\\ &=\sum_{\begin{subarray}{c}I\subset\{1,\ldots,N\}\\ |I|=r\end{subarray}}\prod_{\begin{subarray}{c}j\in I\\ k\notin I\end{subarray}}\frac{\sinh(\pi(x_{j}-x_{k}-ib)/a_{\delta})}{\sinh(\pi(x_{j}-x_{k})/a_{\delta})}\prod_{l\in I}\exp(-ia_{-\delta}\partial_{x_{l}}),\end{split}$ (2.6)
@@
$A_{r,\delta}J_{N}(x,y)=S_{r}(e^{\pi y_{1}/a_{\delta}},\ldots,e^{\pi y_{N}/a_{\delta}})J_{N}(x,y),$ (2.7)
with the elementary symmetric functions
$S_{r}(t)\equiv\sum_{1\leq j_{1}<\cdots j_{r}\leq N}t_{j_{1}}\cdots t_{j_{r}},$ (2.8)
@@
$J_{N}(b;x,y)=\frac{\exp\left(\frac{2\pi i}{a_{+}a_{-}}y_{N}\sum_{j=1}^{N}x_{j}\right)}{(N-1)!}\\ \cdot\int_{{\mathbb{R}}^{N-1}}dzW_{N-1}(b;z){\cal S}^{\sharp}_{N}(b;x,y)J_{N-1}(b;z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N})).$ (2.10)
@@
Since $J_{N}$ is constructed almost entirely from $G(z)$, it is easily seen to have the invariance property
@@ 2.2 Baker–Akhiezer functions
$\nu=\sum_{1\leq j<k\leq N}\left(\frac{m}{2}-l_{jk}\right)(e_{j}-e_{k}),\ \ \ l_{jk}=0,\ldots,m.$ (2.17)
Suppose, in addition, that $\psi_{N}$ satisfies the vanishing condition
$\psi_{N}(x,y+s(e_{j}-e_{k})/2)-\psi_{N}(x,y-s(e_{j}-e_{k})/2)=0,\ \ \ q^{2(y_{j}-y_{k})}=1,$ (2.18)
@@
$\psi_{N,\rho_{N}}(x)=\prod_{1\leq j<k\leq N}\prod_{j=1}^{m}[j+x_{k}-x_{j}],\ \ \ [z]\equiv q^{z}-q^{-z},$ (2.19)
yields a self-dual BA function, in the sense that $\psi_{N}(x,y)=\psi_{N}(y,x)$. Here, we have used the standard notation $\rho_{N}\equiv\rho_{N}(m)$ with
$\rho_{N}(m)\equiv\frac{m}{2}\sum_{1\leq j<k\leq N}(e_{j}-e_{k})=\frac{m}{2}\sum_{j=1}^{N}(N-2j+1)e_{j},$ (2.20)
@@ 3 Baker–Akhiezer specialisation
Taking $m\in\mathbb{Z}_{+}$ and letting
@@
which is easily verified by a direct computation; and if we restrict attention to $|q|=1$ using the parametrisation
$q=e^{-i\pi/a},\ \ \ a\in{\mathbb{R}},$ (3.3)
then it is readily seen that
$D^{r}_{N}(q^{2ix};q^{2},q^{2m})=\sum_{\begin{subarray}{c}I\subset\{1,\ldots,N\}\\ |I|=r\end{subarray}}\prod_{\begin{subarray}{c}j\in I\\ k\notin I\end{subarray}}\frac{\sinh\big(\frac{\pi}{a}(x_{j}-x_{k}-im)\big)}{\sinh\big(\frac{\pi}{a}(x_{j}-x_{k})\big)}\prod_{l\in I}\exp(-i\partial_{x_{l}}).$ (3.4)
@@
$\Phi_{N}(a,m;x,y)=C_{N}\frac{\sum_{\sigma\in S_{N}}(-)^{\sigma}\psi_{N}(\exp(-i\pi/a),m-1;i\sigma(x),iy)}{\prod_{1\leq j<k\leq N}2\sinh\big(\frac{\pi}{a}(y_{k}-y_{j})\big)\prod_{n=-m+1}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{k}-x_{j}+in)\big)}$ (3.5)
with the constant $C_{N}\equiv C_{N}(a,m)$ given by
$C_{N}=(2i^{m})^{N(N-1)/2}\left(\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a}\right)^{(N-1)(N-2)/2}.$ (3.6)
Before proceeding to our proof of the theorem, we note that the weight function $W_{N}$ (2.4) and kernel function ${\cal S}^{\sharp}_{N}$ (2.9) reduce to elementary functions when $b=m\in\mathbb{Z}_{+}$. More precisely, from the difference equation (2.5) satisfied by $G(z)$, we see that
$W_{N}(1,a,m;x)=\prod_{1\leq j\neq k\leq N}\prod_{n=0}^{m-1}2\sinh\left(\frac{\pi}{a}(x_{j}-x_{k}-in)\right)$ (3.7)
and
We note that, up to an overall numerical factor, $W_{N}(1,a,m;x)$ equals the specialisation $\Delta(q^{2ix};q^{2},q^{2m})$ of Macdonald’s weight function $\Delta$, while ${\cal S}^{\sharp}_{N}(1,a,m;x,y)$ is obtained by a similar specialisation of his product function $\Pi$ after removing a factor depending only on $\sum_{j=1}^{N}(x_{j}-y_{j})$; cf. Eqs. (9.2) and (2.5), respectively, in Section VI of [Mac95]. Furthermore, (2.5) and the formula $G(1,a;i(1-a)/2)=1/\sqrt{a}$ (see Eq. (3.38) in [Rui97]) entail that
$\Phi_{N}(x,y)=\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a(N-1)!}e^{\frac{2\pi i}{a}y_{N}(x_{1}+\cdots+x_{N})}\\ \cdot\int_{\mathbb{R}^{N-1}}dz\frac{\Phi_{N-1}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))\prod_{1\leq j\neq k\leq N-1}\prod_{n=0}^{m-1}2\sinh\left(\frac{\pi}{a}(z_{j}-z_{k}-in)\right)}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\cosh\big(\frac{\pi}{a}(x_{j}-z_{k}+i(m-2n-1)/2)\big)}$ (3.10)
with $\Phi_{1}(x,y)=\exp\left(\frac{2\pi}{a}xy\right)$.
As it turns out, the above integrand is a meromorphic function whose poles are due solely to the $\cosh$-factors in the denominator. The first main observation in our proof of Thm. 3.1 is that $\Phi_{N}(x,y)\prod_{1\leq j<k\leq N}2\sinh(\pi(y_{j}-y_{k}))$ can be rewritten as the symmetrisation of a function $\varphi_{N}(x,y)$, given recursively by an iterated residue formula.
@@
Let $q=e^{-i\pi/a}$. Then, under the above assumptions on $m$ and $a$, the BA-function $\psi_{N}$ is obtained from $\psi_{N-1}$ by the iterated residue formula
$\psi_{N}(ix,iy)\\ =\left(\frac{\prod_{n=1}^{m}\sin(\pi n/a)}{2ai^{m+1}}\right)^{N-1}\frac{W_{N}(1,a,m+1;z)}{\prod_{1\leq j<k\leq N}2\sinh\left(\frac{\pi}{a}(x_{k}-x_{j})\right)}e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}\\ \cdot\int_{\underline{\gamma}}dz\frac{\psi_{N-1}(iz,i(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))\prod_{1\leq j<k\leq N-1}2\sinh\left(\frac{\pi}{a}(z_{k}-z_{j})\right)}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n)\big)\big)},$ (3.12)
@@
In the two subsections below, we provide our proofs of Thm. 3.1 and Prop. 3.2. To begin with, we establish the $N=2$ cases of the results. These we then use as the basis for proofs by induction on $N$.
@@ 3.1 The $N=2$ case
Assuming $m\in\mathbb{Z}_{+}$ and a>m1a>m-1, we have
@@
$\varphi_{2,l}(x)=\frac{2a(-i)^{m}}{\prod_{n\neq l}2\sin\big(\frac{\pi}{a}(n-l)\big)}\frac{1}{\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+i(n-l))\big)}.$ (3.18)
Next, we show that $\varphi_{2}$ has vanishing properties similar to those of $\psi_{2}$ given by (2.18) with $N=2$.
@@
$\varphi_{2}\left(x,y+is(e_{1}-e_{2})/2\right)-\varphi_{2}\left(x,y-is(e_{1}-e_{2})/2\right)=0,\ \ \ e^{\frac{2\pi}{a}(y_{1}-y_{2})}=1.$ (3.19)
@@
$(-i)^{m}e^{\frac{2\pi i}{a}(y_{1}-y_{2})x_{1}}\sum_{l=0}^{m-1}\frac{2a}{\prod_{n\neq l}\sin(\frac{\pi}{a}(n-l))}\\ \cdot\frac{\sinh(\frac{\pi s}{a}(x_{2}-x_{1}-i(m-2l-1))}{\prod_{n=0}^{m-1}2\sinh(\frac{\pi}{a}(x_{2}-x_{1}+i(n-l))}e^{-\frac{\pi}{a}(m-2l-1)(y_{1}-y_{2})}.$ (3.21)
@@
$\sum_{l=0}^{m-1}\frac{2a}{\prod_{n\neq l}\sin(\frac{\pi}{a}(n-l))}\frac{\sinh(\frac{\pi s}{a}(x_{2}-x_{1}-i(m-2l-1))}{\prod_{n=0}^{m-1}2\sinh(\frac{\pi}{a}(x_{2}-x_{1}+i(n-l))}=0.$ (3.22)
@@
$\varphi_{2}(x,y)=\frac{2a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\frac{\psi_{N}(\exp(-i\pi/a),m-1;i\sigma(x),iy)}{\prod_{n=-m+1}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+in)\big)}.$ (3.27)

### v2

@@ 1 Introduction
where $(T_{q,x_{l}}f)(x_{1},\ldots,x_{k},\ldots,x_{N})=f(x_{1},\ldots,qx_{l},\ldots,x_{N})$. (The precise relationship can be gleaned from (2.6) and (3.9).) They act on the space of symmetric polynomials in $N$ variables, on which they are simultaneously diagonalised by the symmetric $GL_{N}$ type Macdonald polynomials; see e.g. [Mac95, Nou23, Sto21]. For parameter values of the form $t=q^{m}$, $m\in{\mathbb{Z}}$, Etingof and Styrkas [ES98] and Chalykh [Cha02] constructed and studied non-symmetric joint eigenfunctions of Baker–Akhiezer (BA) type and obtained, in particular, a generalized Weyl character formula for the Macdonald polynomials, first conjectured by Felder and Varchenko [FV97], where the BA-function replaces the exponential function; see also Chalykh and Etingof [CE13] for further related results and references.
@@
For the coupling values $g=m\hbar$ under consideration, our main result provides an expansion of symmetric joint $S_{r}$-eigenfunctions in terms of eigenfunctions that are ‘asymptotically free’ deep inside Weyl chambers $x_{\sigma(1)}>x_{\sigma(2)}>\cdots>x_{\sigma(N)}$, $\sigma\in S_{N}$. In the case of the Macdonald operators and generic parameter values, such asymptotically free solutions, often referred to as (q-)Harisch–Chandra series, have been studied in detail by Letzter and Stokman [LS08], van Meer and Stokman [vMS10] as well as Noumi and Shiraishi [NS12]; and Stokman [Sto14] derived a corresponding ($c$-function) expansion of Cherednik’s basic hypergeometric function associated to root systems. We should also note that Bullimore et al. [BKK15] showed that the partition function of the superconformal field theory $T[U(N)]$, as introduced by Gaiotto and Witten [GW09], displays the very same recursive structure as the joint $S_{r}$-eigenfunctions from [HR14]; and, identifying integration with summing residues at all (infinitely many) poles above the real axis, they obtained an expression in terms of series of an asympototically free type. Furthermore, in the $N=2$ case, di Francesco et al. [FKKSS24] recently obtained a number of interesting results on the $S_{r}$-eigenfunctions, including a similar residue computation where the relevant series are identified as (q-)Harish–Chandra series. To establish the precise connection between these expansion results and the recursive construction in [HR14] is, I believe, an interesting problem to which I plan to return to elsewhere. In the $BC_{1}$ case, the analogous connection between integral representations and series expansions was established by van de Bult et al. [vdBRS07].
The remainder of the paper is structured as follows: In Section 2, we briefly recall definitions and key properties pertaining to joint eigenfunctions of the Ruijsenaars operators (1.1) and BA-(eigen)functions for the Macdonald operators (1.3); and, in Section 3, we give the precise formulations and proofs of our results.
@@ 2.1 Joint eigenfunctions
see loc. cit. Prop. III.1 and Prop. III.2. (Unless needed, we suppress dependence of $G$ and functions constructed from $G$ on the parameters $a_{\pm}$.) Indeed, using these difference equations, it is readily verified that
$\begin{split}A_{r,\delta}(b;x)&\equiv W_{N}(b;x)^{-1/2}H_{r,\delta}(b;x)W_{N}(b;x)^{1/2}\\ &=\sum_{\begin{subarray}{c}I\subset\{1,\ldots,N\}\\ |I|=r\end{subarray}}\prod_{\begin{subarray}{c}j\in I\\ k\notin I\end{subarray}}\frac{\sinh(\pi(x_{j}-x_{k}-ib)/a_{\delta})}{\sinh(\pi(x_{j}-x_{k})/a_{\delta})}\prod_{l\in I}\exp(-ia_{-\delta}\partial_{x_{l}}),\end{split}$ (2.6)
@@
$A_{r,\delta}(b;x)J_{N}(b;x,y)=J_{N}(b;x,y)\sum_{1\leq j_{1}<\cdots<j_{r}\leq N}e^{2\pi(y_{j_{1}}+\cdots+y_{j_{r}})/a_{\delta}},$ (2.7)
@@
$J_{N}(b;x,y)=\frac{\exp\left(\frac{2\pi i}{a_{+}a_{-}}y_{N}\sum_{j=1}^{N}x_{j}\right)}{(N-1)!}\\ \cdot\int_{{\mathbb{R}}^{N-1}}dzW_{N-1}(b;z){\cal S}^{\sharp}_{N}(b;x,z)J_{N-1}(b;z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N})).$ (2.9)
@@
Since $J_{N}$ is constructed almost entirely from $G(z)$, it is easily seen to have the invariance properties
$J_{N}(a_{+},a_{-},b;x,y)=J_{N}(a_{-},a_{+},b;x,y),$ (2.11)
@@ 2.2 Baker–Akhiezer functions
$\nu=\sum_{1\leq j<k\leq N}\left(\frac{m}{2}-l_{jk}\right)(e_{j}-e_{k}),\ \ \ l_{jk}=0,\ldots,m,$ (2.17)
with $\{e_{1},\ldots,e_{N}\}$ the standard ON-basis in $\mathbb{R}^{N}$. Suppose, in addition, that $\psi_{N}$ satisfies the vanishing condition
$\psi_{N}(x,y+s(e_{j}-e_{k})/2)-\psi_{N}(x,y-s(e_{j}-e_{k})/2)=0\ \ \text{when}\ \ q^{2(y_{j}-y_{k})}=1,$ (2.18)
@@
$\psi_{N,\rho_{N}}(x)=\prod_{1\leq j<k\leq N}\prod_{n=1}^{m}[n+x_{k}-x_{j}],\ \ \ [z]\equiv q^{z}-q^{-z},$ (2.19)
yields a self-dual BA function, in the sense that $\psi_{N}(x,y)=\psi_{N}(y,x)$. Here, we have used the standard notation
$\rho_{N}\equiv\rho_{N}(m)=\frac{m}{2}\sum_{1\leq j<k\leq N}(e_{j}-e_{k})=\frac{m}{2}\sum_{j=1}^{N}(N-2j+1)e_{j},$ (2.20)
@@ 3 Baker–Akhiezer specialisation
When $b=m\in\mathbb{Z}_{+}$, the weight function $W_{N}$ (2.4) and kernel function ${\cal S}^{\sharp}_{N}$ (2.8) reduce to elementary functions. More precisely, from the difference equation (2.5) satisfied by $G(z)$, we see that
$W_{N}(1,a,m;x)=\prod_{1\leq j<k\leq N}\sinh\left(\frac{\pi}{a}(x_{k}-x_{j})\right)\prod_{n=-m+1}^{m-1}2\sinh\left(\frac{\pi}{a}(x_{k}-x_{j}+in)\right)$ (3.1)
and
Up to an overall numerical factor, $W_{N}(1,a,m;x)$ equals the specialisation $\Delta(q^{2ix};q^{2},q^{2m})$ of Macdonald’s weight function $\Delta$, while ${\cal S}^{\sharp}_{N}(1,a,m;x,y)$ is obtained by a similar specialisation of his product function $\Pi$ after removing a factor depending only on $\sum_{j=1}^{N}(x_{j}-y_{j})$; cf. Eqs. (9.2) and (2.5), respectively, in Section VI of [Mac95]. Furthermore, (2.5) and the formula $G(1,a;i(1-a)/2)=1/\sqrt{a}$ (see Eq. (3.38) in [Rui97]) entail that
$\Phi_{N}(m;x,y)=\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a(N-1)!}e^{\frac{2\pi i}{a}y_{N}(x_{1}+\cdots+x_{N})}\\ \cdot\int_{\mathbb{R}^{N-1}}dz\prod_{1\leq j<k\leq N-1}\sinh\left(\frac{\pi}{a}(z_{k}-z_{j})\right)\prod_{n=-m+1}^{m-1}2\sinh\left(\frac{\pi}{a}(z_{k}-z_{j}+in)\right)\\ \cdot\frac{\Phi_{N-1}(m;z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\cosh\big(\frac{\pi}{a}(x_{j}-z_{k}+i(m-2n-1)/2)\big)}$ (3.4)
with $\Phi_{1}(x,y)=\exp\left(\frac{2\pi}{a}xy\right)$. We note that this is effectively a basic hypergeometric integral, since the weight and kernel functions can be expressed in terms of the product
$\prod_{n=0}^{m-1}(q^{n+z}-q^{-n-z})=(-1)^{m}q^{-m(m-1)/2-mz}\frac{(q^{2z};q^{2})_{\infty}}{(q^{2(m+z)};q^{2})_{\infty}},$ (3.5)
with
$q=e^{-i\pi/a},\ \ \ a\in\mathbb{R}.$ (3.6)
To be precise, while the (modified) $q$-Gamma function $(q^{2z};q^{2})_{\infty}$ is only well-defined for $|q|<1$, the above ratio clearly extends to $|q|=1$. For accounts of (multivariable) basic hypergeometric integrals, see e.g. the overview by Koornwinder and Stokman [KS21] and the survey by Schlosser [Sch21].
Upon setting $(a_{+},a_{-},b)=(1,a,m)$, the $a$-antiperiodicity of the functions involved entails that the eigenvalue equations (2.7), which are are also satisfied by $\Phi_{N}$, become trivial for $\delta=+$. As we now show, the remaining equations, corresponding to $\delta=$-, are closely related to the Macdonald eigenvalue equations (2.21), of which the BA-functions are solutions.
Letting
@@
which is easily verified by a direct computation; and with the parametrisation (3.6) in effect it is readily seen that
$q^{mr(1-N)}D^{r}_{N}(q^{2ix};q^{2},q^{2m})=\sum_{\begin{subarray}{c}I\subset\{1,\ldots,N\}\\ |I|=r\end{subarray}}\prod_{\begin{subarray}{c}j\in I\\ k\notin I\end{subarray}}\frac{\sinh\big(\frac{\pi}{a}(x_{j}-x_{k}-im)\big)}{\sinh\big(\frac{\pi}{a}(x_{j}-x_{k})\big)}\prod_{l\in I}\exp(-i\partial_{x_{l}}).$ (3.9)
@@
[…]
$\mathcal{\delta}_{N}(i\sigma(x);m-1)=(-)^{\sigma}\mathcal{\delta}_{N}(ix;m-1),\ \ \ \sigma\in S_{N},$ (3.10)
which is clear from (3.7). Moreover, given that $\Phi_{N}(a,m;x,y)$ is manifestly symmetric in $x$ (cf. (2.9) and (2.14)), the coefficients in the linear combination should be of the form $c(y)(-)^{\sigma}$, with $c$ independent of $\sigma$. In this section, we substantiate this expectation by proving the following theorem.
@@
$\Phi_{N}(a,m;x,y)\\ =C_{N}(a,m)\frac{\sum_{\sigma\in S_{N}}(-)^{\sigma}\psi_{N}(e^{-i\pi/a},m-1;i\sigma(x),iy)}{\prod_{1\leq j<k\leq N}2\sinh\big(\frac{\pi}{a}(y_{k}-y_{j})\big)\prod_{n=-m+1}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{k}-x_{j}+in)\big)}$ (3.11)
with the constant
$C_{N}(a,m)=(i^{m})^{N(N-1)/2}\left(\frac{a}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\right)^{(N-1)(N-2)/2}.$ (3.12)
The first main observation in our proof of Thm. 3.1, as detailed below, is that the product function $\Phi_{N}(a,m;x,y)\prod_{1\leq j<k\leq N}2\sinh(\pi(y_{j}-y_{k}))$ can be rewritten as the symmetrisation of a function $\varphi_{N}(a,m;x,y)$ that is given recursively by an iterated residue formula.
@@
Under the above assumptions on $m$ and $a$, the BA-function $\psi_{N}$ is obtained from $\psi_{N-1}$ by the iterated residue formula
$\psi_{N}(e^{-i\pi/a},m;ix,iy)\\ =\left(\frac{\prod_{n=1}^{m}\sin(\pi n/a)}{a(-i)^{m+1}}\right)^{N-1}\frac{W_{N}(1,a,m+1;z)}{\prod_{1\leq j<k\leq N}2\sinh\left(\frac{\pi}{a}(x_{k}-x_{j})\right)}e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}\\ \cdot\int_{\underline{\gamma}}dz\frac{\psi_{N-1}(e^{-i\pi/a},m;iz,i(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))\prod_{1\leq j<k\leq N-1}2\sinh\left(\frac{\pi}{a}(z_{k}-z_{j})\right)}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n)\big)\big)},$ (3.14)
@@
In the two subsections below, we provide our proofs of Thm. 3.1 and Prop. 3.2. To begin with, we establish the $N=2$ cases of the results. These we then use as the basis for proofs by induction on $N$. To shorten some of the formulae involved, we typically suppress dependence on the parameters $a$ and $m$.
@@ 3.1 The $N=2$ case
Assuming that $m\in\mathbb{Z}_{+}$ and a>m1a>m-1, we have
@@
$\varphi_{2,l}(x)=\frac{a(-i)^{m}}{\prod_{n\neq l}2\sin\big(\frac{\pi}{a}(n-l)\big)}\frac{1}{\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+i(n-l))\big)}.$ (3.20)
We note that the above series essentially amounts to a special (terminating) case of the basic hypergeometric series
$_{2}\phi_{1}\left(\,\begin{matrix}a,\,b\\ \,c\end{matrix}\,\Big|\,q,z\right)=\sum_{l=0}^{\infty}\frac{(a:q)_{l}(b:q)_{l}}{(c;q)_{l}(q;q)_{l}}z^{l},$ (3.21)
where the $q$-shifted factorial is given by $(a;q)_{0}=1$ and $(a;q)_{l}=(1-a)\cdots(1-aq^{l-1})$ when $l>0$. To verify this claims, it suffices to observe that
$\begin{split}\frac{1}{\prod_{n\neq l}2\sin\left(\frac{\pi}{a}(n-l)\right)}&=\frac{1}{\prod_{n=1}^{m-1}2\sin\left(\frac{\pi}{a}n\right)}\prod_{n=0}^{l-1}\frac{2\sin\left(\frac{\pi}{a}(m-1-n)\right)}{2\sin\left(\frac{\pi}{a}(-n-1)\right)}\\ &=\frac{t_{m}^{l}}{\prod_{n=1}^{m-1}2\sin\left(\frac{\pi}{a}n\right)}\frac{(q^{2}/t_{m}^{2};q^{2})_{l}}{(q^{2};q^{2})_{l}}\end{split}$ (3.22)
and similarly that
$\frac{1}{\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+i(n-l))\big)}\\ =t_{m}^{l}\frac{1}{\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+in)\big)}\frac{(q^{2}X_{2}/t_{m}^{2}X_{1};q^{2})_{l}}{(q^{2}X_{2}/X_{1};q^{2})_{l}},$ (3.23)
where we have used the parametrisation (3.6) for $q$ and introduced
$t_{m}\equiv e^{-\frac{i\pi}{a}m}=q^{m}$ (3.24)
as well as
$X_{j}\equiv e^{\frac{2\pi}{a}x_{j}}=q^{i2x_{j}},\ \ Y_{j}\equiv e^{\frac{2\pi}{a}y_{j}}=q^{i2y_{j}},\ \ \ j=1,2.$ (3.25)
Indeed, it is clear from the above computations that
$\begin{split}\varphi_{2}(x,y)&=\frac{a(-i)^{m}\exp\left(\frac{2\pi i}{a}(x-i\rho_{2}(m-1),y)\right)}{\prod_{n=1}^{m-1}2\sin(\pi n/a)\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+in)\big)}\\ &\quad\cdot{}_{2}\phi_{1}\left(\,\begin{matrix}q^{2}/t_{m}^{2},\,q^{2}X_{2}/t_{m}^{2}X_{1}\\ \,q^{2}X_{2}/X_{1}\end{matrix}\,\Big|\,q^{2},t_{m}^{2}Y_{2}/Y_{1}\right).\end{split}$ (3.26)
Starting from the $q$-difference eigenvalue equation satisfied by $_{2}\phi_{1}$, it would now be a simple exercise to show that $\varphi_{2}$ is an eigenfunction of the $N=2$ Macdonald operator $D_{2}^{1}$ for suitable parameter values. One could also invoke the duality relation (2.15) and linear independence of $\varphi_{2}(\sigma(x),y)$, $\sigma\in S_{2}$, for generic $x\in\mathbb{R}^{2}$; see Lemma 3.8 below. Either way, one could then use a uniqueness result such as Prop. 4.1 in [Nou23] for ‘asymptotically free’ eigenfunctions of $D_{2}^{1}$ to arrive at the connection between $\varphi_{2}$ and $\psi_{2}$. We note that such uniqueness results for arbitrary $N$ can be found in the paper of Letzter and Stokman [LS08] (in a general root system setting) as well as that of Noumi and Shiraishi [NS12].
Here, we rely instead on the defining properties of $\psi_{2}$ and show that $\varphi_{2}$ has vanishing properties similar to those given by (2.18) with $N=2$.
@@
$\varphi_{2}\left(x,y+is(e_{1}-e_{2})/2\right)-\varphi_{2}\left(x,y-is(e_{1}-e_{2})/2\right)=0\ \ \text{when}\ \ e^{\frac{2\pi}{a}(y_{1}-y_{2})}=1.$ (3.27)
@@
$(-i)^{m}e^{\frac{2\pi i}{a}(y_{1}-y_{2})x_{1}}\sum_{l=0}^{m-1}\frac{a}{\prod_{n\neq l}2\sin(\frac{\pi}{a}(n-l))}\\ \cdot\frac{\sinh(\frac{\pi s}{a}(x_{2}-x_{1}-i(m-2l-1))}{\prod_{n=0}^{m-1}2\sinh(\frac{\pi}{a}(x_{2}-x_{1}+i(n-l))}e^{\frac{\pi}{a}(m-2l-1)(y_{1}-y_{2})}.$ (3.29)
@@
$\sum_{l=0}^{m-1}\frac{a}{\prod_{n\neq l}2\sin(\frac{\pi}{a}(n-l))}\frac{\sinh(\frac{\pi s}{a}(x_{2}-x_{1}+i(m-2l-1))}{\prod_{n=0}^{m-1}2\sinh(\frac{\pi}{a}(x_{2}-x_{1}+i(n-l))}=0.$ (3.30)
@@
$\varphi_{2}(x,y)=\frac{2a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\frac{\psi_{2}(\exp(-i\pi/a),m-1;i\sigma(x),iy)}{\prod_{n=-m+1}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+in)\big)}.$ (3.35)